# Non-Markovianity of subsystem dynamics in isolated quantum many-body systems

### v1

@@ Abstract
It is believed that an isolated quantum many-body system far away from equilibrium should try to attain equilibrium via a mechanism whereby any given subsystem acts as an open quantum system that is coupled to an environment which is the complementary part of the full system and undergoing a complicated equilibration process such that all the subsystems in the long-time limit attain equilibrium states compatible with the global equilibrium state. This picture begs the question whether the dynamics of any given subsystem is Markovian (monotonic loss of information and memory) or non-Markovian. In this work, by numerically probing the dynamical behaviour of the distance between temporally-separated subsystem states, we reveal the telltale signatures and other associated features of quantum (non-)Markovianity of the dynamics of small subsystems of an isolated quantum spin system in one dimension (the Ising spin chain) brought far from equilibrium by quantum quenches. Additionally, remarkably systematic behaviour is seen in the dynamics of the eigenvalues of the reduced density matrices of the subsystems. These features strongly depend on the direction of quenching in the parameter space, with paramagnetic-to-ferromagnetic quenches offering considerably stronger signatures of subsystem non-Markovianity, for which we offer heuristic arguments.
@@
Introduction— It is expected that a closed quantum many-body system when brought far out of equilibrium by quenching one of its parameters will after a sufficiently long-time equilibriate to a state as dictated by the post-quench parameters. The actual process of equilibration can be rather complex because during this process the system will explore the set of accessible states (decided by the extant conserved quantities) in the Hilbert space during its Hamiltonian dynamics, and the complexity of the trajectories in the quantum state space depends crucially on the initial state and its relation with the Hamiltonian driving its evolution. For generic (i.e., non-integrable) systems with only the total energy being a conserved quantity, the long-time physics should then be described by usual statistical mechanics in the microcanonical picture, which essentially encapsulates the notion of ergodicity. One then says that the system has thermalized, i.e., the long-time (thermal) state is the Gibbs density matrix as described by the Gibbs ensemble, and the eigenstate thermalization hypothesis (ETH) describes well the behaviour of observables in this regime [1, 2, 3, 4]. Exceptions to this paradigm within the class of non-integrable quantum many-body systems exist of course [5, 6, 7, 8]. For integrable quantum many-body systems, i.e., those with an extensive number of conserved quantities, the long-time physics is instead described by a so-called generalized Gibbs ensemble picture wherein the equilibrated state is a generalized Gibbs density matrix [9, 10], and one says that these systems do not thermalize in the usual sense (but possibly in a generalized sense [11]) and constitute a strong violation of ergodicity and the ETH.
@@
With this line of thinking, and employing the notion of distances or divergences between quantum states of which many measures exist in (quantum) information theory [12, 13, 14], our work investigates the following basic question : how does the distance between temporally separated quantum states of a given subsystem evolve in time and what underlying characteristics of the non-equilibrium dynamics are thereby revealed. Independently, the viewpoint of regarding a given subsystem as an open quantum system necessitates asking whether its dynamics is Markovian (memory-less, and monotonic loss of information) or not [15, 16, 17, 18]. Both these questions are intimately related to each other, and we believe this classification of subsystem dynamics as Markovian or non-Markovian is a concrete addition to the existing list of characteristics of the non-equilibrium dynamics of closed quantum many-body systems.
@@
Distances between quantum states— We choose the trace distance as our primary measure of distance between quantum states in this article (results are qualitatively unchanged if other measures are chosen). Given two density matrices $\rho$ and $\sigma$, the trace distance (TD) between them is defined as $T_{d}(\rho,\sigma)\!=\!\frac{1}{2}\sum|\lambda_{i}|$, where $\{\lambda_{i}\}$ is the set of eigenvalues of $(\rho-\sigma)$, with the index $i\!=\!1,2,..,\operatorname{rank}(\rho-\sigma)$ The trace distance takes values $\in[0,1]$. Let $\{p^{\downarrow}\}$ and $\{q^{\downarrow}\}$ be the vector of eigenvalues of $\rho$ and $\sigma$ respectively, arranged in a descending order. The counterpart of trace distance for probability distributions, known as the total variation distance (TVD), is defined as $V_{d}(p^{\downarrow},q^{\downarrow})\!=\!\frac{1}{2}\sum_{i}|p^{\downarrow}_{i}-q^{\downarrow}_{i}|$, and it also takes values $\in[0,1]$.
@@
Divisibility, information backflow and non-Markovianity— A fundamental and powerful property of any distance measures is that they are non-increasing (or "contractive") under CPTP maps [12, 13, 14, 19]. That is, given a CPTP operation $\Lambda:\mathcal{D(\mathcal{H})}\!\rightarrow\!\mathcal{D(\mathcal{H})}$, one has $T_{d}(\Lambda(\rho),\Lambda(\sigma))\leq T_{d}(\rho,\sigma)$ for the trace distance. In other words, any (physically realizable) quantum operation can not increase the (information about the) distance between (and hence, a quantifier of "distinguishability") of two quantum states.
@@
Consider now a dynamically evolving quantum system. Let $\Lambda_{t,0}$ denote the family of CPTP maps that evolves the subsystem from initial time $t=0$ to time $t$ for all $t>0$, these will be referred to as dynamical maps. This temporal evolution process is said to be (CP-)divisible if one has $\Lambda_{t,0}=\Lambda_{t,s}\Lambda_{s,0}$ for any $s\in(0,t)$, where $\Lambda_{t,s}$ is also a CPTP map evolving the system from time $s$ to $t$, otherwise if $\Lambda_{t,s}$ is not CPTP, the evolution is termed (CP-)indivisible (note that it is not to be taken for granted nor is it obvious or generic that the intermediate map $\Lambda_{t,s}$ has to be CPTP). It then follows that in a divisible process, the aforementioned monotonicities are respected at every step of the process, and so ultimately one has $T_{d}(\Lambda_{t,0}(\rho),\Lambda_{t,0}(\sigma))\leq T_{d}(\rho,\sigma)$ for any $t>0$. The distance between two quantum states continues to decrease all along the temporal dynamics, and such a dynamical process is termed Markovian, and any violation of this condition can be taken as a signature of non-Markovianity [20, 15, 16]. Thus, an CP-indivisible process is non-Markovian though the converse does not necessarily hold true [21]. An alternative approach is to view the upholding of these monotonicities as monotonic loss of information ("memory") about the distinguishable features of a pair of quantum systems coupled separately to the same environment (which is stable by itself and is ideally not affected by its interactions with the non-equilibrium quantum systems to which it is coupled, which is the essence of the Born-Markov approach to open quantum systems [22, 23]), whereby the environment acts to drive the systems to the final equilibrium states and thereby decreasing any distinguishable features (absorbing the information about such distinguishabilities) between them over time; such a process is Markovian by this token [16]. Consequently, any violation of these monotonicities signal backflows of information from the environment to the systems (i.e., the environment somehow retained the dynamical memory of the past states of the systems; the system-environment coupling is beyond the Born-Markov regime and the environment is in fact itself affected by the systems it is coupled to and unable to quickly shed information about them) and this is a defining signature of non-Markovianity [24, 25]. These two viewpoints are not necessarily equivalent to each other in general nor is it obvious that they should be, however for invertible CPTP maps [[Invertibilityisnottobeconfusedwithreversibility, andinvertiblemapsarenotgenerallyCPTPeither.OnlyforaunitaryCPTPmapisitsinversealsoCPTPandconsequentlyinvertibilityisthesameasreversibility, seefore.g.][.]Chruscinski2022a] (that is, $\Lambda_{a,b}^{-1}$ exists, meaning none of the eigenvalues of $\Lambda_{a,b}$ is zero; such is the case for us as explicitly checked numerically) these two viewpoints are understood to be equivalent [27, 28, 29, 18]. To be sure, our numerical results are really only measuring the backflows of information, and not actually establishing if all the intermediate $\Lambda_{t,s}$ are CPTP or not and thereby establishing (in)divisibility from first principles, which is a separate issue in itself and is being addressed in a parallel work. It is only thanks to these maps being invertible that we talk of both these notions of non-Markovianity interchangeably in this work.
@@
Consider then a subsystem (not necessarily contiguous) of length $l$ spins, represented by a density matrix $\rho_{t}^{(\ell)}$ at time $t$, of a non-equilibrium quantum many-body system that is evolving under its own Hamiltonian ($t=0$ denotes the starting time). We shall be dealing exclusively with translationally invariant systems in this work, thus it does not matter where exactly this $-\ell$ subsystem is located within the full system as long as it is located in the bulk away from the boundaries (in case of open boundary conditions). We probe the following basic questions in this work : how the distance between two temporally-separated states of a subsystem behaves as the system evolves in time, and what characteristics of the non-equilibrium dynamics is revealed by this behaviour, and if there are any invariants or qualitative changes under varying temporal separation of considered subsystem states ? As such, our main quantities of interest are $T_{d}(\rho_{t+\delta}^{\ell},\rho_{t}^{\ell})$, where $\delta$ is the temporal separation. Let $\Lambda_{[t]}$ be the dynamical CPTP map that evolves the subsystem by time $t$, i.e., $\rho_{t+a}^{\ell}\!=\!\Lambda_{[t]}[\rho_{a}^{\ell}]$, then by previous discussion, if the dynamics is divisible, i.e., $\Lambda_{[t]}\!=\!(\Lambda_{[1]})^{t}$ (more generally, in terms of the previous discussion, $\Lambda_{[t]}\!=\!\Lambda_{t+a,a}$ for any reference time stamp $a$), then a non-increasing behaviour of $T_{d}(\rho_{t+\delta}^{\ell},\rho_{t}^{\ell})$ must be observed for any $\delta$. Consequently, any violation of this monotonically non-increasing behaviour of these two quantities is a signature of indivisibility as well as "backflow" of information (about the distinguishability between quantum states) from the environment back to the subsystem in question and thus, of non-Markovianity, and a degree of non-Markovianity may then be defined by the magnitude of such violations.
@@
Model and methods— We demonstrate our results on the paradigmatic one-dimensional Ising spin chain in the presence of both transverse and longitudinal fields with open boundary conditions,
$\mathcal{H}_{I}=-J\sum_{j=1}^{N-1}\sigma^{z}_{j}\sigma_{j+1}^{z}-h_{x}\sum_{j=1}^{N}\sigma_{j}^{x}-h_{z}\sum_{j=1}^{N}\sigma_{j}^{z}\text{ .}$
@@
When $h_{z}\!=\!0$, the model is integrable, with the Ising critical point at $J\!=\!h_{x}$ separating the symmetry-broken ferromagnetic/antiferromagnetic phases from the paramagnetic phase [30]. When $h_{z}\!\neq\!0$, the model is non-integrable and in the symmetry-broken phases (ferromagnetic or anti-ferromagnetic) exhibits confinement between kink-antikink excitations [31, 32, 33, 34] and consequently slow thermalization after a quench to this non-integrable regime [35, 36, 37, 38, 39, 40, 41, 42, 43, 44].
@@
For our simulations we use the matrix product state (MPS) formalism and second-order time-evolving block decimation (TEBD2) algorithm for simulating the real-time dynamics [45, 46, 47]. Our simulations were performed with the ITensors library in Julia [48, 49]. Most results shown here were obtained with Trotter time-steps $\tau\!=\!0.01$ seconds for our TEBD2 simulations (thus, our numerical errors are $\mathcal{O}(10^{-4})$), and we have checked that the results were unchanged with a smaller time-step of $\tau\!=\!0.002$ seconds. The total system size was fixed at $N\!=\!200$ sites, and results were independent of the other system sizes used for verification purposes. The MPS cutoffs were set at $10^{-9}$, and maximum bond-dimensions were set at $50$. The ground states were obtained with the density matrix renormalization group (DMRG) algorithm [46], also implemented with the same cutoffs as mentioned above.
In this work we have focused on investigating small subsystems comprising of a few spins embedded in the (translationally-invariant) bulk of a large spin chain. This is partly due to the fact that the computational costs of constructing reduced density matrices scale exponentially in the subsystem sizes, and in fact already for subsystems of only five spins this was evident to us in practice. This also partly served the purpose of going systematically in investigating the dynamics of subsystems starting up from the smallest possible one. We will see below that already for a four-spin subsystem, the non-Markovianity of its dynamics is quite insignificant, so our focus on small subsystems gets incentivized a posteriori. Perhaps most importantly, in contemporary ultracold atomic experiments, local site-resolved studies of quantum states has been made feasible and put to practice (see e.g. [50, 51]), which motivates putting the microscope on very small subsystems also theoretically.
@@
Figure 1: Paramagnetic-to-ferromagnetic quench in mixed-field Ising spin chain: Trace distances (TD) as functions of time for various small subsystem RDMs with $\delta=1$ sec. (left), and $\delta=2$ sec. (right). The highly non-monotonic behaviour signifies strong non-Markovianity.
@@
Figure 2: Degree of subsystem non-Markovianity for TD distances vs $\delta$ in the mixed-field Ising spin chain. Upper Row - for various small subsystems in paramagnetic to ferromagnetic quench (left), and ferromagnetic to paramagnetic quench (right). Lower Row - for 4-RDMs with varying longitudinal field (left) and transverse field (right).

### v2

@@ Abstract
It is believed that an isolated and far-from-equilibrium quantum many-body system should try to attain equilibrium via a mechanism whereby any given subsystem acts as an open quantum system that is coupled to an environment, which is the complementary part of the full system, and undergoes a complicated equilibration process such that all the subsystems in the long-time limit attain equilibrium states compatible with the global equilibrium state. This picture begs the question whether the dynamics of any given subsystem is Markovian (monotonic loss of information and memory) or non-Markovian. In this work, by numerically probing the dynamical behaviour of the quantum distances between temporally-separated states of small subsystems, we reveal the telltale signatures of (non-)Markovianity of the dynamics of subsystems of an isolated quantum spin system brought in the far-from-equilibrium regime, exemplified with the mixed-field Ising spin chain quenched between parameter regimes deep inside its magnetically ordered and disordered regimes. Additionally, remarkably systematic behaviour is seen in a measure of classical distances between the quantum states of the considered subsystems. These features strongly depend on the direction of quenching in the parameter space, with paramagnetic-to-ferromagnetic quenches offering considerably stronger signatures of subsystem non-Markovianity, for which we offer heuristic arguments.
I Introduction
@@
It is expected that a closed quantum many-body system when brought far out of equilibrium by quenching one of its parameters will, after a sufficiently long-time, equilibriate to a state as dictated by the post-quench parameters. The actual process of equilibration can be rather complex because during this process the system will explore the set of accessible states (decided by the extant conserved quantities) in the Hilbert space during its Hamiltonian dynamics, and the complexity of the trajectories in the quantum state space depends crucially on the initial state and its relation with the Hamiltonian driving its evolution. For generic (i.e., non-integrable) systems with only the total energy being a conserved quantity, the long-time physics is described by usual statistical mechanics in the microcanonical picture, which essentially encapsulates the notion of ergodicity. One then says that the system has thermalized, i.e., the long-time (thermal) state is the Gibbs density matrix as described by the Gibbs ensemble, and the eigenstate thermalization hypothesis (ETH) describes well the behaviour of observables in this regime [1, 2, 3, 4]. Exceptions to this paradigm within the class of non-integrable quantum many-body systems exist of course [5, 6, 7, 8]. For integrable quantum many-body systems, i.e., those with an extensive number of conserved quantities, the long-time physics is instead described by a so-called generalized Gibbs ensemble picture wherein the equilibrated state is a generalized Gibbs density matrix [9, 10], and one says that these systems do not thermalize in the usual sense (but possibly in a generalized sense [11]) and constitute a strong violation of ergodicity and the ETH.
@@
With this line of thinking, and employing the notion of distances or divergences between quantum states of which many measures exist in (quantum) information theory [12, 13, 14], this work investigates the following basic question : how does the distance between temporally separated quantum states of a given subsystem evolve in time and what underlying characteristics of the non-equilibrium dynamics are thereby revealed. Independently, the viewpoint of regarding a given subsystem as an open quantum system necessitates asking whether its dynamics is Markovian (memory-less, and monotonic loss of information) or not [15, 16, 17, 18, 19]. Both these questions are intimately related to each other, and we believe this classification of subsystem dynamics as Markovian or non-Markovian is a concrete addition to the existing list of characteristics of the non-equilibrium dynamics of closed quantum many-body systems (see also [20]).
In this work we have focused on investigating small subsystems comprising of a few spins embedded in the (translationally-invariant) bulk of a large spin chain. This is partly due to the fact that the computational costs of constructing reduced density matrices (RDM) scale exponentially in the subsystem sizes, and in fact already for subsystems of only five spins this was evident to us in practice. This also partly served the purpose of going systematically in investigating the dynamics of subsystems starting up from the smallest possible one. We will see later that already for a four-spin subsystem, the non-Markovianity of its dynamics is quite insignificant, so our focus on small subsystems gets incentivized a posteriori. Perhaps most importantly, in contemporary ultracold atomic experiments, local site-resolved studies of quantum states has been made feasible and put to practice (see e.g. [21, 22, 23]), which motivates putting the microscope on very small subsystems also theoretically. In fact, given that experimental access and control is often limited to small subsystems and not the whole of large quantum many-body systems, probing and understanding the various attributes of the dynamics of small subsystems (as experimentally accessible open systems surrounded by the inaccessible environments) is thus also a pertinent theoretical task.
This article is organized as follows. We introduce our notations and some preliminary notions in Section II, Section III introduces and discusses the concept and a common measure of information backflow, and Markovianity and its lack thereof, as well as how to quantify them in the context of isolated quantum many-body dynamics. In Section IV, we present our results for the mixed-field Ising spin chain, using the trace distance measure (defined in Eq.(1)) as our figure of merit, with the main result being that the paramagnetic-to-ferromagnetic quenches induce significant information backflow and non-Markovianity of subsystem dynamics, while the oppositely-directed quenches exhibit essentially Markovian dynamics of small subsystems. In addition, notably systematic behaviour in the dynamical behaviour of a corresponding "classical" distance measure (defined in Eq.(2)) is also revealed and discussed. Section V provides an extended discussion of several aspects of our work and related matters, and Section VI provides concluding remarks.
II Preliminaries
@@
Distances between quantum states— We choose the trace distance as our primary measure of distance between quantum states in this article (results are qualitatively unchanged if other measures are chosen). Our choice for this measure is motivated by simplicity and ease of computability, as the trace distance measure between quantum states is free from any potential numerical issues that may arise with other measures involving fractional powers or logarithms of sufficiently large density matrix operators. Given two density matrices $\rho$ and $\sigma$, the trace distance (TD) between them is defined as,
$T_{d}(\rho,\sigma)=\frac{1}{2}\sum|\lambda_{i}|\text{ ,}$ (1)
@@
where $\{\lambda_{i}\}$ is the set of eigenvalues of $(\rho-\sigma)$, with the index $i\!=\!1,2,..,\operatorname{rank}(\rho-\sigma)$ The trace distance takes values $\in[0,1]$. Let $\{p^{\downarrow}\}$ and $\{q^{\downarrow}\}$ be the vector of eigenvalues of $\rho$ and $\sigma$ respectively, arranged in a descending order. The counterpart of trace distance for probability distributions, known as the total variation distance (TVD), is defined as,
$V_{d}(p^{\downarrow},q^{\downarrow})=\frac{1}{2}\sum_{i}|p^{\downarrow}_{i}-q^{\downarrow}_{i}|\text{ ,}$ (2)
and it also takes values $\in[0,1]$. This can be thought of as the classical counterpart to the trace distance, and encodes in some sense a notion of classical distances between the respective density matrices.
III Assessing (non-)Markovianity
@@
A fundamental and powerful property of any distance measures is that they are non-increasing (or "contractive") under CPTP maps [12, 13, 14, 24]. That is, given a CPTP operation $\Lambda:\mathcal{D(\mathcal{H})}\!\rightarrow\!\mathcal{D(\mathcal{H})}$, one has, for the trace distance,
$T_{d}(\Lambda(\rho),\Lambda(\sigma))\leq T_{d}(\rho,\sigma)\text{ .}$ (3)
In other words, any (physically realizable) quantum operation can not increase the (information about the) distance between (and hence, a quantifier of "distinguishability") of two quantum states.
Consider now a dynamically evolving quantum system. Let $\Lambda_{t,0}$ denote the family of CPTP maps that evolves the subsystem from initial time $t=0$ to time $t$ for all $t>0$, these will be referred to as dynamical maps. This temporal evolution process is said to be (CP-)divisible if one has,
$\Lambda_{t,0}=\Lambda_{t,s}\Lambda_{s,0}\hskip 8.5359pt\forall s\in(0,t)\text{,}$ (4)
where $\Lambda_{t,s}$ is also a CPTP map evolving the system from time $s$ to $t$, otherwise if $\Lambda_{t,s}$ is not CPTP, the evolution is termed (CP-)indivisible (note that it is not to be taken for granted nor is it obvious or generic that the intermediate map $\Lambda_{t,s}$ has to be CPTP) [25]. It then follows that in a divisible process, the aforementioned monotonicities are respected at every step of the process, and so ultimately one has,
$T_{d}(\Lambda_{t,0}(\rho),\Lambda_{t,0}(\sigma))\leq T_{d}(\rho,\sigma)\text{ ,}$ (5)
@@
for any $t>0$. The distance between two quantum states continues to decrease all along the temporal dynamics, and such a dynamical process is termed Markovian, and any violation of this condition can be taken as a signature of non-Markovianity [26, 15, 16]. Thus, an CP-indivisible process is non-Markovian though the converse does not necessarily hold true [27]. An alternative approach is to view the upholding of these monotonicities as reflecting the monotonic loss of information ("memory") about the distinguishable features of a pair of quantum systems coupled separately to the same environment, which is stable by itself and is not affected by its interactions with the non-equilibrium quantum systems to which it is coupled (which is the essence of the Born-Markov approach to open quantum systems [28, 29]). Such an environment acts to drive the systems to the final equilibrium states and thereby decreasing any distinguishable features (absorbing the information about such distinguishabilities) between them over time; such a process is Markovian by this token [16]. Consequently, any violation of these monotonicities signal backflows of information from the environment to the systems (i.e., the environment somehow retained the dynamical memory of the past states of the systems; the system-environment coupling is beyond the Born-Markov regime and the environment is in fact itself affected by the systems it is coupled to and unable to quickly shed information about them) and this is a defining signature of non-Markovianity [30, 31]. These two viewpoints are not necessarily equivalent to each other in general nor is it obvious that they should be, however for invertible CPTP maps (that is, $\Lambda_{a,b}^{-1}$ exists, meaning none of the eigenvalues of $\Lambda_{a,b}$ is zero; such is the case for us as explicitly checked numerically) these two viewpoints are understood to be equivalent [32, 33, 34, 18]. It is important to remark here that invertibility is not to be confused with reversibility, and invertible maps are not generally CPTP either. Only for a unitary CPTP map is its inverse also CPTP and consequently invertibility is the same as reversibility, see e.g. [18] for extended discussions on this subtle point. Moreover, for certain classes of dynamical maps, invertibility has been argued to be necessary for Markovianity [35], though clearly it can not be sufficient. To be sure, our numerical results are really only measuring the backflows of information, and not actually establishing if all the intermediate $\Lambda_{t,s}$ are CPTP or not and thereby establishing (in)divisibility from first principles, which is a separate issue in itself and is being addressed in a parallel work. It is only thanks to these maps being invertible that we talk of both these notions of non-Markovianity interchangeably in this work.
@@
Consider then a subsystem (not necessarily contiguous) of length $l$ spins, represented by a density matrix $\rho_{t}^{(\ell)}$ at time $t$, of a non-equilibrium quantum many-body system that is evolving under its own Hamiltonian ($t=0$ denotes the starting time). We shall be dealing exclusively with translationally invariant systems in this work, thus it does not matter where exactly this $-\ell$ subsystem is located within the full system as long as it is located in the bulk away from the boundaries (in case of open boundary conditions). We probe the following basic questions in this work : how the distance between two temporally-separated states of a subsystem behaves as the system evolves in time, and what characteristics of the non-equilibrium dynamics is revealed by this behaviour, and if there are any invariants or qualitative changes under varying temporal separation of considered subsystem states ? As such, our main quantities of interest are $T_{d}(\rho_{t+\delta}^{\ell},\rho_{t}^{\ell})$, where $\delta$ is the temporal separation. Let $\Lambda_{(t)}$ be the dynamical CPTP map that evolves the subsystem by time $t$, i.e.,
$\rho_{t+a}^{\ell}=\Lambda_{(t)}[\rho_{a}^{\ell}]\text{ .}$ (6)
Then by previous discussion, if the dynamics is divisible, i.e., $\Lambda_{(t)}\!=\!(\Lambda_{(1)})^{t}$ (more generally, in terms of the previous discussion, $\Lambda_{(t)}\!=\!\Lambda_{t+a,a}$ for any reference time stamp $a$), then a non-increasing behaviour of $T_{d}(\rho_{t+\delta}^{\ell},\rho_{t}^{\ell})$ and consequently no information backflow must be observed for any $\delta$. Consequently, any violation of this monotonically non-increasing behaviour of these two quantities is a signature of indivisibility as well as a backflow of information (about the distinguishability between quantum states) from the environment back to the subsystem in question and thus, of non-Markovianity, and a degree of non-Markovianity may then be defined by the cumulative magnitude of such violations. Speaking in terms of discrete times as in numerical simulations (continuous-time generalization is obvious) with a time-step $\tau$, let the discrete "slope" of the aforementioned trace distance for a given $\delta$ be denoted as,
$\alpha(t,\delta)=\frac{1}{\tau}\bigg{(}T_{d}(\rho_{t+\tau+\delta}^{\ell},\rho_{t+\tau}^{\ell})-T_{d}(\rho_{t+\delta}^{\ell},\rho_{t}^{\ell})\bigg{)}\text{ .}$ (7)
This quantity is always negative for Markovian dynamics. Thus, a degree of non-Markovianity measuring the cumulative magnitude of revivals (increases) of the trace distance $T_{d}(\rho_{t+\delta}^{\ell},\rho_{t}^{\ell})$ can be defined as,
[…]
Note that this differs from the degree of non-Markovianity used in [30, 31] where a maximization over pairs of initial states has been made due to their interests in defining this degree for a given quantum dynamical process acting on various initial states. Our interest however is slightly different, in that we are concerned with a fixed initial state for a given class of quenching and wish to quantify the non-Markovianity arising out of quenching it by parameters lying in a different phase of the underlying system, and then comparing this degree for different quenching parameters within the same post-quench parameter regime. That is, for example, in paramagnetic-to-ferromagnetic class of quenches, the initial paramagnetic phase is fixed and a non-Markovianity degree is calculated for each ferromagnetic quenching parameters, and then these degrees are compared with each other for different ferromagnetic quenching parameters. Later in Sec.IV, we shall also consider the dynamical behaviour of the trace variation distance Eq.(2) between the descendingly-ordered eigenvalues of the temporally separated subsystem density matrices, and a quantity measuring the cumulative magnitude of revivals of the TVD measure can be similarly defined. Let the corresponding slope for the TVD measure for a given $\delta$ be,
$\alpha_{1}(t,\delta)=\frac{1}{\tau}\bigg{(}V_{d}(q_{t+\tau+\delta}^{\ell},q_{t+\tau}^{\ell})-V_{d}(q_{t+\delta}^{\ell},q_{t}^{\ell})\bigg{)}\text{ ,}$ (9)
where $q_{t}^{\ell}$ denotes the eigenvalues, arranged in descending order, of the density matrix $\rho_{t}^{\ell}$. The degree of TVD revivals can then be defined as,
$\mathcal{D}_{1}(\delta)=\sum_{t}\alpha_{1}(t,\delta)\hskip 8.5359pt\forall t\hskip 7.11317pt\text{s.t.}\hskip 7.11317pt\alpha_{1}(t,\delta)>0\text{ .}$ (10)
IV Results and discussion for the mixed-field Ising spin chain
@@
We demonstrate our results on the paradigmatic one-dimensional Ising spin chain in the presence of both transverse and longitudinal fields with open boundary conditions,
$\mathcal{H}_{I}=-J\sum_{j=1}^{N-1}\sigma^{z}_{j}\sigma_{j+1}^{z}-h_{x}\sum_{j=1}^{N}\sigma_{j}^{x}-h_{z}\sum_{j=1}^{N}\sigma_{j}^{z}\text{ .}$ (11)
@@
When $h_{z}\!=\!0$, the model is integrable, with the Ising critical point at $J\!=\!h_{x}$ separating the symmetry-broken ferromagnetic/antiferromagnetic phases from the paramagnetic phase [36]. When $h_{z}\!\neq\!0$, the model is non-integrable and in the symmetry-broken phases (ferromagnetic or anti-ferromagnetic) exhibits confinement between kink-antikink excitations [37, 38, 39, 40] and consequently slow thermalization (and other features associated with it) after quenching to this non-integrable regime [41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51].
@@
For our simulations we use the matrix product state (MPS) formalism and second-order time-evolving block decimation (TEBD2) algorithm for simulating the real-time dynamics [52, 53, 54]. Our simulations were performed with the ITensors library in Julia [55, 56]. Most results shown here were obtained with Trotter time-steps $\tau\!=\!0.01$ for our TEBD2 simulations (thus, our numerical errors are $\mathcal{O}(10^{-4})$), and we have checked that the results were unchanged with a smaller time-step of $\tau\!=\!0.002$. The total system size was fixed at $N\!=\!200$ sites, and results were independent of the other system sizes used for verification purposes. The MPS cutoffs were set at $10^{-9}$, and maximum bond-dimensions were set at $50$. The ground states were obtained with the density matrix renormalization group (DMRG) algorithm [53], also implemented with the same cutoffs as mentioned above. The unit of time is $J^{-1}$ for paramagnetic-to-ferromagnetic quenches, and $h_{x}^{-1}$ for ferromagnetic-to-paramagnetic quenches.
@@
Figure 1: Paramagnetic-to-ferromagnetic quench $(J,h_{x},h_{z})\!=\!(0.2,1,0)\rightarrow(1,0.1,0.5)$ : Trace distances (TD), Eq.(1) (between $\delta-$separated RDMs of small subsystems), plotted on the y-axes vs time for various small subsystem RDMs with $\delta\!=\!1$ (left), and $\delta\!=\!2$. (right). In the labels, $k-$RDM refers to subsystem size, $k\!=\!\{1,2,3,4\}$. The highly non-monotonic behaviour signifies strong non-Markovianity.
@@
Figure 2: Degree of subsystem non-Markovianity for TD distances $\mathcal{D}(\delta)$, Eq.(8), on the y-axes vs $\delta$. Upper Row - For various small subsystems in paramagnetic to ferromagnetic quench $(J,h_{x},h_{z})\!=\!(0.2,1,0)\rightarrow(1,0.1,0.5)$ (left), and ferromagnetic to paramagnetic quench $(J,h_{x},h_{z})\!=\!(1,0.1,0.5)\rightarrow(0.2,1,0)$ (right). Note that this degree of subsystem non-Markovianity in the latter quench are smaller than in the former quench by a factor of about 50 or more. Lower Row - For 4-RDMs with varying longitudinal field $h_{z}$ (left) and transverse field $h_{x}$ (right) in the case of paramagnetic-to-ferromagnetic quenches.